# TOMATOES: Topology and Material Optimization for Latent Heat Thermal Energy Storage Devices

### v1

@@ 3.2 Design Representation and Material Interpolation
We employ standard element-wise pseudo-densities [47] to represent the design. In particular, the pseudo-density of element $e$ is $\gamma_{e}\in[0,1]$, where a value of zero indicates the element is composed of the HCM, and a value of one indicates the element is composed of the PCM. Observe that although the pseudo-densities may assume any continuous value in [0,1], we require a binary design at convergence. To this end, we use a material interpolation scheme that penalizes intermediate densities. In particular, the effective thermal conductivity is obtained using the standard Solid Isotropic Material with Penalization (SIMP) scheme [11]:
@@
Note that Equation˜8 assumes the latent heat of the HCM is zero. This modeling approximation is justified given the HCM serves only to enhance conduction and the operating temperatures are assumed well below its melting point. Similarly, the melting temperature is fixed to that of the PCM and is not interpolated.
Additionally, the pseudo-densities are regularized using a standard density filter to ensure mesh-independence and preclude checker-boarding. A subsequent threshold projection is applied to the filtered field to obtain a binary design. The specifics of these filtering and projection methods are detailed in [48].
@@ 3.3 Differentiable Material Representation
Building upon the density-based design representation, we now detail the representation of discrete material databases in a continuous and differentiable form, amenable for gradient-based optimization. To this end, we employ variational autoencoders (VAEs). VAEs are generative models that have been applied for compression, semi-supervised learning, and interpolation [49]. In design optimization, they have been used to generate parametric microstructures [50], photonic devices [30], and robust designs [51]. Following the approach of [7], we leverage the ability of VAEs to map a discrete database onto a continuous, differentiable latent space that can be integrated into a gradient-based optimization framework.
@@
The HCM and PCM VAEs are trained on databases containing $n_{H}=56$ and $n_{P}=140$ materials, respectively. The training process minimizes a loss function composed of a mean squared error reconstruction term and a Kullback–Leibler (KL) divergence regularizer. This regularizer encourages the posterior distribution of the latent space to approximate a standard normal distribution, $\mathcal{N}(0,\mathbf{I})$ [14]. The total loss function is given by Equation˜9, where the terms $\bm{\hat{\Psi}}$ and $\bm{\Psi}$ collectively represents the input and reconstructed material properties respectively, and the weighting factor $\beta$ is set to $10^{-7}$.
@@ 3.4 Finite Element Analysis
For the temporal discretization, we employ an implicit, first-order backward Euler scheme. This yields the discretized system of equations to compute the temperature field at time step $n+1$ as Equation˜10.
@@
The effective mass density $\rho_{eff}$ and apparent heat capacity $\breve{c}_{eff}$ once again follow the interpolations as discussed in Section˜3.2. Furthermore, observe that the apparent heat capacity is a function of temperature, making the system of equations nonlinear. The resulting system of nonlinear residual equations is solved at each time step using a modified Newton-Raphson iterative scheme, as detailed in [52]. To enhance the robustness and accuracy of the numerical solution, a Gradient Galerkin Least Squares (GGLS) stabilization technique is incorporated [53, 54].
To ensure computational performance, our FEA solver, developed within the JAX ecosystem [55], is integrated with the high-performance PETSc library [56] for the solution of the linear systems of equations at each Newton iteration. Critically, the entire computational framework is constructed to be end-to-end differentiable. This enables the efficient and accurate computation of design sensitivities via automatic differentiation; discussed further in Section˜3.6.
@@ 3.5 Optimization
Latent Space Constraint: Observe that while the optimizer continuously varies the latent coordinates $\vec{z}_{H}$ and $\vec{z}_{P}$ during optimization, the materials encoded from the database occupy discrete points in the latent space. In other words, the optimizer is allowed to explore materials outside the database during optimization. However, at convergence, we expect to select materials exclusively from the library. To enforce this, we impose a constraint on the distances between the latent coordinates of selected material instances and those of the database materials. With $\bm{\Delta}_{H}$ and $\bm{\Delta}_{P}$ being the pairwise distances between the latent coordinate of the selected material instance and database materials for the HCM and PCM respectively:
@@
Note that to facilitate gradient-based optimization, we use the LogSumExp approximations of the min and max functions [57]. Furthermore, we use a continuation scheme on the allowed maximum distance $\epsilon^{*}$ according to the strategy detailed in Section˜4.
@@
The constrained minimization problem Equation˜19 is then transformed into an unconstrained loss function minimization, using the log-barrier scheme [58]. Specifically, the loss function is defined as
@@
and $J_{0}$ is the initial objective. The constraint penalty parameter $\tau$ is updated using a continuation scheme, making the enforcement of the constraint stricter as the optimization progresses (see Section˜4). The gradient-based Adam optimizer [59] is used to minimize Equation˜20. The details for obtaining the gradients are discussed in the next section.
@@ 3.6 Sensitivity Analysis
A critical step in gradient-based topology optimization is the computation of sensitivities, which are the derivatives of the objective and constraint functions with respect to the design variables [28]. The sensitivity analysis for this work is particularly involved because it must handle transient, nonlinear thermal physics and a computational pipeline that integrates FEA with a neural network (in particular, the material decoders).
To address this complexity, we construct an end-to-end differentiable pipeline using the automatic differentiation (AD) capabilities of the JAX framework [55]. This allows us to avoid the laborious and error-prone process of manually deriving complex sensitivity expressions. By implementing the entire forward analysis within JAX—from the VAE-based material mapping to the FEA solver—the framework computes the required derivatives to machine precision using reverse-mode AD.
Additionally, we emphasize on two challenges inherent to our physics simulation. The first arises from the use of iterative schemes, such as the Newton-Raphson method, to solve the nonlinear governing equations. A naive application of AD would unroll the derivative calculations through every solver iteration, a process that is both computationally expensive and memory-intensive. To overcome this inefficiency, we apply the Implicit Function Theorem (IFT) [60, 28], which enables the direct computation of derivatives from the final converged solution, thereby bypassing the need to backpropagate through the iterative solution history.
The second challenge stems from the transient nature of the simulation, which creates a significant memory bottleneck for the adjoint sensitivity method. An adjoint analysis requires access to the thermal state from all previous time steps to compute the gradient at the current step, and storing this entire state history is often infeasible for long simulations. We mitigate this issue by employing a checkpointing scheme [61, 62]. This technique stores the system’s state at select time steps, and during the reverse-time adjoint solve, states between these checkpoints are recomputed on the fly. This method significantly reduces memory requirements at the cost of a moderate increase in computation time, making the analysis of long-duration transient problems feasible.
@@ 3.7.2 Co-Geometry-Material Optimization
    Design Update: The objective value, constraint values, and their sensitivities are passed to the ADAM optimizer [59], which then computes an updated set of design variables.
This process (Steps 3–9) is repeated iteratively until the change in design variables between iterations falls below a specified tolerance, or a maximum number of iterations is reached. The optimal material choice and its distribution are obtained as outputs of the optimization.
@@ 4 Numerical Experiments
In this section, we present several experiments to demonstrate the proposed framework. All experiments were conducted on a MacBook M2 Air, using the JAX library [55] in Python. Unless otherwise specified, the default parameters for all numerical examples are as follows:
@@
    Filter: A convolution type filter with a radius of 0.03m0.03\;m is applied to the material pseudo-densities to mitigate checkerboarding and ensure mesh-independence [63].
@@
    Thermal Model: The inner boundary is maintained at a temperature $T_{d}$ of 273K273\;K (Equation˜1) while the other boundaries are considered as insulated. The entire domain has a uniform initial temperature of $T_{I}=400\;K$ (Equation˜1). The reference temperature in Equation˜14 is taken to be $T_{d}$. The transient analysis is conducted for 60 time steps with a step size of $8000\;\text{seconds}$ (133 hours).
@@
    Numerical Solvers: The nonlinear system of equations at each time step is solved using a modified Newton-Raphson method [52] with a convergence tolerance of $10^{-7}$ for the residual norm. The linear system is solved using the PETSc [64] solver.
@@
    Optimizer: The optimization is performed using the ADAM optimizer [59] with a learning rate of $5\times 10^{-2}$.
@@
    Continuation Schemes: To promote convergence and smoother optimizations, we adopt continuation schemes ([63]) for the following parameters:
@@ 4.1 Validation of Hypothesis
We begin by establishing the advantage of employing TO for the geometric design, with pre-selected materials. We assume commonly used materials: Aluminum as the HCM and RT-25 Paraffin as the PCM. We compare the energy discharge performance of five designs as in Figure˜7: a domain filled entirely with PCM, traditional two-fin, three-fin, and four-fin configurations, and the TO design. The pure PCM design serves as a baseline. While it possesses the maximum theoretical latent energy capacity, its poor thermal conductivity limits heat extraction. This results in a low net discharged energy of only $5\text{ MJ}$ within the discharge period. In contrast, the traditional fin designs, with an allocated $20\%$ of the total volume to the HCM, demonstrate significantly improved performance. Although the inclusion of the HCM reduces the total latent heat capacity, it creates crucial conductive pathways that facilitate more effective heat extraction, thereby increasing the total energy discharged.
@@
Figure 7: Designs with pure PCM, 2-fins, 3-fins, 4-fins, geometry optimization and energy discharge plot.
Having demonstrated the importance of geometric optimization, we now compare the sequential design strategy against our proposed concurrent co-design strategy. The results of this comparison are illustrated in Figure˜8 and Figure˜9. In the sequential optimization strategy, we use the optimized geometry obtained in the previous step (using pre-selected materials, Aluminum and RT-25 Paraffin). Then, keeping the geometry and PCM fixed, we optimize the HCM choice, which results in the selection of Boron Nitride and an improved energy discharge of $229.5\text{ MJ}$. Next, using this selected HCM and geometry, we select the optimal PCM. This sequential optimization yields the optimal PCM choice as X90 PlusICE, and a discharge of $268.5\text{ MJ}$.
@@
Figure 8: Sequential and co-design of geometry and material.
Figure 9: Comparison of energy discharge curves for different design strategies.
@@ 4.2 Pareto Tradeoff and Convergence
A common consideration in engineering design is the trade-off between performance and cost. To demonstrate this, we consider the effect of varying the maximum allowable cost for the HCM. Keeping other parameters including the maximum discharge time, and operating temperatures constant, we perform the co-optimization to obtain the Pareto front as illustrated in Figure˜10. As expected, the discharged energy increases with the allowable cost.
@@
Finally, Figure˜11 shows the convergence history for the design corresponding to the $\mathdollar 300$ cost constraint. The objective and constraint values show stable convergence. This optimization was completed in 15 minutes and 32 seconds; similar convergence characteristics were observed for the other experiments.
@@ 4.3 Variation in Operating Temperature
Figure˜12 shows the final optimized designs. The framework selected a different PCM for each operating temperature range. For the $350\text{ K}$ and $450\text{ K}$ cases, it chose X80 PlusICE and X90 PlusICE, respectively. For the high-temperature $550\text{ K}$ case, it chose Bismuth. Each selected PCM has a melting point that falls within its target operating range. This indicates the framework selects appropriate materials. The optimal fin geometry also changes with temperature, once again highlighting the interdependence of geometry and material.

### v2

@@ 3.2 Design Representation and Material Interpolation
We employ standard element-wise pseudo-densities [sigmund200199] to represent the design. In particular, the pseudo-density of element $e$ is $\gamma_{e}\in[0,1]$, where a value of zero indicates the element is composed of the HCM, and a value of one indicates the element is composed of the PCM. Observe that although the pseudo-densities may assume any continuous value in [0,1], we require a binary design at convergence. To this end, we use a material interpolation scheme that penalizes intermediate densities. In particular, the effective thermal conductivity is obtained using the standard Solid Isotropic Material with Penalization (SIMP) scheme [pizzolato2017LHTES]:
@@
Note that Equation˜8 assumes the latent heat of the HCM is zero. This modeling approximation is justified, given that the HCM serves only to enhance conduction, and the operating temperatures are assumed well below its melting point. Similarly, the melting temperature is fixed to that of the PCM and is not interpolated.
Additionally, the pseudo-densities are regularized using a standard density filter to ensure mesh-independence and preclude checker-boarding. A subsequent threshold projection is applied to the filtered field to obtain a binary design. The specifics of these filtering and projection methods are detailed in [wu2017infill].
@@ 3.3 Differentiable Material Representation
Building upon the density-based design representation, we now detail the representation of discrete material databases in a continuous and differentiable form, amenable to gradient-based optimization. To this end, we employ variational autoencoders (VAEs). VAEs are generative models that have been applied for compression, semi-supervised learning, and interpolation [kingma2019VAE]. In design optimization, they have been used to generate parametric microstructures [padhy2024tomas], photonic devices [padhy2024photos], and robust designs [gladstone2021robustTOVAE]. Following the approach of [chandrasekhar2022integrating], we leverage VAEs’ ability to map a discrete database to a continuous, differentiable latent space that can be integrated into a gradient-based optimization framework.
@@
The HCM and PCM VAEs are trained on databases containing $n_{H}=56$ and $n_{P}=140$ materials, respectively. The training process minimizes a loss function composed of a mean-squared-error reconstruction term and a Kullback–Leibler (KL) divergence regularizer. This regularizer encourages the posterior distribution of the latent space to approximate a standard normal distribution, $\mathcal{N}(0,\mathbf{I})$ [kingma2013auto]. The total loss function is given by Equation˜9, where the terms $\bm{\hat{\Psi}}$ and $\bm{\Psi}$ collectively represent the input and reconstructed material properties, respectively, and the weighting factor $\beta$ is set to $10^{-7}$.
@@ 3.4 Finite Element Analysis
For temporal discretization, we employ an implicit first-order backward-Euler scheme. This yields a discretized system of equations for computing the temperature field at time step $n+1$ as in Equation˜10.
@@
The effective mass density $\rho_{eff}$ and apparent heat capacity $\breve{c}_{eff}$ once again follow the interpolations as discussed in Section˜3.2. Furthermore, observe that the apparent heat capacity is a function of temperature, making the system of equations nonlinear. The resulting system of nonlinear residual equations is solved at each time step using a modified Newton-Raphson iterative scheme, as detailed in [alexandersen2023detailed]. To enhance the robustness and accuracy of the numerical solution, a Gradient Galerkin Least Squares (GGLS) stabilization technique is incorporated [pizzolato2018thesis, franca1989galerkin].
To ensure computational performance, our FEA solver, developed within the JAX ecosystem [jax2018github], integrates with the high-performance PETSc library [petsc] to solve linear systems at each Newton iteration. Critically, the entire computational framework is constructed to be end-to-end differentiable. This enables the efficient and accurate computation of design sensitivities via automatic differentiation; discussed further in Section˜3.6.
@@ 3.5 Optimization
Latent Space Constraint: Observe that while the optimizer continuously varies the latent coordinates $\vec{z}_{H}$ and $\vec{z}_{P}$ during optimization, the materials encoded from the database occupy discrete points in the latent space. In other words, the optimizer is allowed to explore materials outside the database during optimization. However, at convergence, we expect to select materials exclusively from the library. To enforce this, we impose a constraint on the distances between the latent coordinates of selected material instances and those of the database materials. With $\bm{\Delta}_{H}$ and $\bm{\Delta}_{P}$ being the pairwise distances between the latent coordinate of the selected material instance and database materials for the HCM and PCM, respectively:
@@
Note that to facilitate gradient-based optimization, we use the LogSumExp approximations of the min and max functions [zhang2023dive]. Furthermore, we use a continuation scheme on the allowed maximum distance $\epsilon^{*}$ according to the strategy detailed in Section˜4.
@@
The constrained minimization problem Equation˜19 is then transformed into an unconstrained loss function minimization, using the log-barrier scheme [kervadec2022logBarrier]. Specifically, the loss function is defined as
@@
and $J_{0}$ is the initial objective. The constraint penalty parameter $\tau$ is updated using a continuation scheme, making the enforcement of the constraint stricter as the optimization progresses (see Section˜4). The gradient-based Adam optimizer [kingma2014adam] is used to minimize Equation˜20. The details for obtaining the gradients are discussed in the next section.
@@ 3.6 Sensitivity Analysis
A critical step in gradient-based topology optimization is the computation of sensitivities, which are the derivatives of the objective and constraint functions with respect to the design variables [padhy2025toflux]. The sensitivity analysis for this work is particularly complex because it must account for transient, nonlinear thermal physics and a computational pipeline that integrates FEA with a neural network (in particular, the material decoders).
To address this complexity, we construct an end-to-end differentiable pipeline using the automatic differentiation (AD) capabilities of the JAX framework [jax2018github]. This allows us to avoid the laborious and error-prone process of manually deriving complex sensitivity expressions. By implementing the entire forward analysis within JAX—from the VAE-based material mapping to the FEA solver—the framework computes the required derivatives to machine precision using reverse-mode AD.
Furthermore, we emphasize two challenges inherent to our physics simulation. The first arises from the use of iterative schemes, such as the Newton-Raphson method, to solve the nonlinear governing equations. A naive application of AD would unroll derivative computation across every solver iteration, a process that is both computationally expensive and memory-intensive. To overcome this inefficiency, we apply the Implicit Function Theorem (IFT) [blondel2022ImplicitDifferentiation, padhy2025toflux], which enables the direct computation of derivatives from the final converged solution, thereby bypassing the need to backpropagate through the iterative solution history.
The second challenge stems from the transient nature of the simulation, which creates a significant memory bottleneck for the adjoint sensitivity method. An adjoint analysis requires access to the thermal state from all previous time steps to compute the gradient at the current step, and storing this entire state history is often infeasible for long simulations. We mitigate this issue by employing a checkpointing scheme [wang2009minimal, james2015topology]. This technique stores the system’s state at select time steps, and during the reverse-time adjoint solve, states between these checkpoints are recomputed on the fly. This method significantly reduces memory requirements at the cost of a moderate increase in computation time, making the analysis of long-duration transient problems feasible.
@@ 3.7.2 Co-Geometry-Material Optimization
    Design Update: The objective value, constraint values, and their sensitivities are passed to the ADAM optimizer [kingma2014adam], which then computes an updated set of design variables.
This process (Steps 3–9) is repeated iteratively until the change in design variables between iterations falls below a specified tolerance, or a maximum number of iterations is reached. The optimal material choice and its distribution are obtained as outputs of the optimization. The flowchart of the optimization is shown in Figure˜7.
Figure 7: The design optimization procedure.
@@ 4 Numerical Experiments
In this section, we present several experiments to demonstrate the proposed framework. All experiments were conducted on a MacBook M2 Air, using the JAX library [jax2018github] in Python. Unless otherwise specified, the default parameters for all numerical examples are as follows:
@@
    Filter: A convolution type filter with a radius of 0.03m0.03\;m is applied to the material pseudo-densities to mitigate checkerboarding and ensure mesh-independence [sigmund1998numericalInstabilites].
@@
    Thermal Model: The inner boundary is maintained at a temperature $T_{d}$ of 273K273\;K (Equation˜1) while the other boundaries are considered as insulated. The entire domain has a uniform initial temperature of $T_{I}=400\;K$ (Equation˜1). The reference temperature in Equation˜14 is taken to be $T_{d}$. The transient analysis is conducted for 60 time steps with a step size of $800\;\text{seconds}$ (13.3 hours).
@@
    Numerical Solvers: The nonlinear system of equations at each time step is solved using a modified Newton-Raphson method [alexandersen2023detailed] with a convergence tolerance of $10^{-7}$ for the residual norm. The linear system is solved using the PETSc [balay2019petsc] solver.
@@
    Optimizer: The optimization is performed using the ADAM optimizer [kingma2014adam] with a learning rate of $5\times 10^{-2}$.
@@
    Continuation Schemes: To promote convergence and smoother optimizations, we adopt continuation schemes ([sigmund1998numericalInstabilites]) for the following parameters:
@@ 4.1 Validation of Hypothesis
We begin by establishing the advantage of employing TO for the geometric design, with pre-selected materials. We assume commonly used materials: Aluminum as the HCM and RT-25 Paraffin as the PCM. We compare the energy discharge performance of five designs as in Figure˜8: a domain filled entirely with PCM, traditional two-fin, three-fin, and four-fin configurations, and the TO design. The pure PCM design serves as a baseline. While it possesses the maximum theoretical latent energy capacity, its poor thermal conductivity limits heat extraction. This results in a low net discharged energy of only $5\text{ MJ}$ within the discharge period. In contrast, the traditional fin designs, with an allocated $20\%$ of the total volume to the HCM, demonstrate significantly improved performance. Although the inclusion of the HCM reduces the total latent heat capacity, it creates crucial conductive pathways that facilitate more effective heat extraction, thereby increasing the total energy discharged.
@@
Figure 8: Performance comparison of baseline and geometry optimized designs using fixed materials. The topology-optimized design (dotted orange line) significantly outperforms standard 2-fin, 3-fin, and 4-fin geometries.
Having demonstrated the importance of geometric optimization, we now compare the sequential design strategy against our proposed concurrent co-design strategy. The results of this comparison are illustrated in Figure˜9 and Figure˜10. In the sequential optimization strategy, we use the optimized geometry obtained in the previous step (using pre-selected materials, Aluminum and RT-25 Paraffin). Then, keeping the geometry and PCM fixed, we optimize the HCM choice, which results in the selection of Boron Nitride and an improved energy discharge of $229.5\text{ MJ}$. Next, using this selected HCM and geometry, we select the optimal PCM. This sequential optimization yields the optimal PCM choice as X90 PlusICE, and a discharge of $268.5\text{ MJ}$.
@@
Figure 9: Comparison of the sequential design strategy (top row) versus the proposed concurrent co-design framework (bottom row). The sequential approach optimizes geometry (A), followed by HCM selection (B), and PCM selection (C) as successive steps. In contrast, the co-design formulation (D) concurrently optimizes topology and material choice.
Figure 10: Temporal evolution of the total stored energy (MJ) for different design strategies. The curves illustrate the stored energy remaining in the system over the discharge period for: (A) geometry optimization only, (B) sequential HCM selection, (C) sequential PCM selection, and (D) concurrent co-design.
@@ 4.2 Pareto Tradeoff and Convergence
A common consideration in engineering design is the trade-off between performance and cost. To demonstrate this, we consider the effect of varying the maximum allowable cost for the HCM. Keeping other parameters including the maximum discharge time, and operating temperatures constant, we perform the co-optimization to obtain the Pareto front as illustrated in Figure˜11. As expected, the discharged energy increases with the allowable cost.
@@
Finally, Figure˜12 shows the convergence history for the design corresponding to the $\mathdollar 300$ cost constraint. The objective and constraint values show stable convergence. This optimization was completed in 15 minutes and 32 seconds; similar convergence characteristics were observed for the other experiments.
@@
4.3 Effect of Initialization
In this experiment, we investigate the influence of the starting design on the optimal design. Given the nonconvex, nonlinear nature of the transient TO problem, we anticipate that different initializations will yield distinct local optima. To study robustness, we performed five independent optimization runs initialized with distinct, randomly sampled points in the material latent space $(\vec{z}_{P},\vec{z}_{H})$. All optimizations were conducted subject to a maximum cost constraint of $\mathdollar 600$.
Figure˜13 illustrates the resulting optimized designs. Observe that while the optimizer converges to diverse geometric topologies and different material combinations, the final performance remains similar (with a maximum deviation of around 4% from the mean). This suggests that, as expected, the loss landscape contains numerous local solutions with similar performance. Crucially, it demonstrates the robustness of our TO framework, which can discover designs with the desired performance regardless of initialization.
Furthermore, a challenge in continuous representation of discrete databases is the physical realizability of the optimized materials, i.e., whether the final solution corresponds to a real material. Table˜1 compares the material properties predicted by decoders at convergence for design (a) against the values for the nearest materials in the database (Graphite and X90 PlusICE). The decoded properties match the actual physical values with a relative error consistent with fig.˜6. This confirms the effectiveness of the latent space constraint (Equation˜18), ensuring the optimizer converges to physically valid materials from the latent space.
Figure 13: Optimized designs for different initializations. The designs (a-e) illustrate local optima found by the framework. The designs (a-e) show the local optima corresponding to the starting latent positions $\vec{z}_{P},\vec{z}_{H}$ listed below each subfigure for PCM and HCM respectively.
Table 1: Comparison of decoded material properties at convergence versus actual database values, verifying physical realizability.
Property HCM (Graphite) PCM (X90 PlusICE)
Decoded Database Decoded Database
Thermal Conductivity ($W/m\cdot K$) 200.73 200.00 0.357 0.360
Density ($kg/m^{3}$) 2197.06 2200.00 1190.54 1200.00
Specific Heat ($J/kg\cdot K$) 709.31 710.00 1539.48 1510.00
Latent Heat ($J/kg$) - - 168,390 170,000
Melting Temp ($K$) - - 362.97 363.00
Cost / Unit (\$) 1.48 1.50 - -
@@ 4.4 Variation in Operating Temperature
Figure˜14 shows the final optimized designs. The framework selected a different PCM for each operating temperature range. For the $350\text{ K}$ and $450\text{ K}$ cases, it chose X80 PlusICE and X90 PlusICE, respectively. For the high-temperature $550\text{ K}$ case, it chose Bismuth. Each selected PCM has a melting point that falls within its target operating range. This indicates the framework selects appropriate materials. The optimal fin geometry also changes with temperature, once again highlighting the interdependence of geometry and material.